OpenStudy (anonymous):

how do i sketch this graph y = x^3 + x^2 - 5x can i get it in detail plz

OpenStudy (anonymous):

Well, by factoring that graph and finding the roots. You can approximate the graph using the end behaviour.

OpenStudy (anonymous):

Since it's degree 3, you know it's an odd function. What's the end behaviour for that?

OpenStudy (anonymous):

Actually i am trying to find the total area enclosed between the curves y= x^3+x^2-5x and y=x^2-x